OpenStudy (toxicsugar22):

A computer store sells approximately 300 of its most popular model per year. The manager would like to minimize her annual inventory cost by ordering the optimal numbers of computers, x, at regular intervals. If she orders x computers in each shipment, the cost of storage willl be 6x dollers, and the cost of reording will be 300/x (15x+10) dollers. The inventory cost is the sum of the storage cost and the reording cost. use the distributive law to simplfy the expression for the rerding cost. Then express the inventory cost C as a function of x

OpenStudy (toxicsugar22):

hi dumbcow

OpenStudy (dumbcow):

hi

OpenStudy (toxicsugar22):

do any of u guys underand this question

OpenStudy (dumbcow):

working on it give me a minute

OpenStudy (dumbcow):

is x(x+15) all on denominator or just the x ?

OpenStudy (anonymous):

the simplified version of ur question is 30x^3 + 20x^2 = 300 correct me if am wrong

OpenStudy (anonymous):

is the cost of reording: ???

OpenStudy (anonymous):

\[{ 300 \over x} (15x+10)\]

OpenStudy (toxicsugar22):

yes math mind got it rightd

OpenStudy (anonymous):

isn't it equal to 6? [300/x(15x + 10) = 6\]

OpenStudy (anonymous):

6x that is

OpenStudy (dumbcow):

\[C = \frac{6(x ^{2}+50x+750)}{x}\]

OpenStudy (anonymous):

Then express the inventory cost C as a function of x:\[C(x) = 6x+{300 \over x}(15x+10) = 6x+{3000 \over x} + 45000\] is that right?

OpenStudy (anonymous):

oh

OpenStudy (dumbcow):

wow i misread the question, i am wrong sorry time for me to leave...haha

OpenStudy (anonymous):

haha. it happens .

OpenStudy (toxicsugar22):

yes math mind got it right

OpenStudy (toxicsugar22):

so it says complete this of valuses for the inventory cost for various reorder size so if x is 10 what isC how would i set that up can u show me

OpenStudy (anonymous):

all yu do i plug 10 for x: \[C(10)=6(10)+{3000 \over 10}+45000 = ?\]

OpenStudy (dumbcow):

\[C=\frac{6(x^2+750x+500)}{x} \]

OpenStudy (toxicsugar22):

ok i got it

OpenStudy (toxicsugar22):

thanks

OpenStudy (toxicsugar22):

How many computers should the manager order in each shipment so as to minimize the inventury cost

OpenStudy (dumbcow):

optimal number of computers to minimize cost is 10sqrt(5)

OpenStudy (dumbcow):

took derivative of C and set equal to zero and solve for x

OpenStudy (anonymous):

what formula are yu using cow?

OpenStudy (toxicsugar22):

math mind what do u think

OpenStudy (dumbcow):

use the quotient rule since it's a fraction

OpenStudy (toxicsugar22):

How many computers should the manager order in each shipment so as to minimize the inventury cost

OpenStudy (toxicsugar22):

well does anyone know

OpenStudy (anonymous):

sorry.. tv distracted me haha.. i got the same thing as cow 10sqrt{5}

OpenStudy (anonymous):

which is 22.36 computers

OpenStudy (toxicsugar22):

ok thanks and how many orders will she make during the year?

OpenStudy (toxicsugar22):

does anyone know this

OpenStudy (anonymous):

if the stores sell aprox. 300 a year and it orders 22.36 a shipment: 300/22.36 = 13.41 i think that's the answer

OpenStudy (toxicsugar22):

ok thanks

OpenStudy (toxicsugar22):

graph this equation and estimate the possible values for can u do that I graphed it and tried various nubers but it is not ight

OpenStudy (toxicsugar22):

can u try

OpenStudy (toxicsugar22):

like if u have a graphing calculator

OpenStudy (anonymous):

oh sorry, that i can't do :/

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

its okay

OpenStudy (toxicsugar22):

do u know a website tgraph it for me

OpenStudy (toxicsugar22):

do u know any website that will graph it for me

OpenStudy (toxicsugar22):

ok thanks